# Causal Attribution via Activation Patching

### 13 Mar 2026 v1

@@ B.2 Input-Based Intervention Baselines
Simple Insertion 55.63±1.24 69.51±1.19 10.63±1.01 43.86±1.10 75.33±0.87 14.93±0.83 49.79±1.20 79.41±1.11 19.59±0.79 33.33±0.92 64.50±0.97 22.16±0.75
Simple Deletion 58.89±1.37 69.75±1.07 10.86±1.20 60.40±1.27 72.38±0.89 11.98±1.14 59.82±1.33 80.18±1.05 20.36±0.93 42.34±0.90 61.13±0.89 18.79±0.88
@@
Table 13: Localization quantitative results for Simple Insertion, Simple Deletion, and CAAP on ImageNet. AUPR1 ($\uparrow$), AUPR0 ($\uparrow$) and PG ($\uparrow$) are reported. Best is bold, second best is underlined in each column.
@@
Simple Insertion 49.28±1.99 71.10±2.18 61.72±4.30 55.32±1.96 75.55±2.09 59.38±4.34 48.13±2.02 72.81±2.16 52.34±4.41 55.59±1.97 79.02±1.99 62.50±4.28
Simple Deletion 40.82±2.07 67.53±2.23 46.09±4.41 40.44±2.10 66.01±2.18 58.59±4.35 39.05±2.09 67.45±2.27 45.31±4.40 47.51±2.50 80.74±1.69 57.81±4.37
@@
To better understand the role of representation-level interventions, we compare CAAP with two simple input-based alternatives: Simple Insertion and Simple Deletion. In Simple Insertion, a selected source patch is inserted directly into a blank target image and the resulting class score is used as the attribution signal. In Simple Deletion, the selected patch is removed from the input image while all other pixels remain unchanged. These procedures operate only at the input level and do not modify internal representations across layers. All methods use the default padding ($3\times 3$ patch region). Tables 12 and 13 report the results on ImageNet across four ViT backbones.
@@
B.3 CNN architectures
Table 14: Faithfulness quantitative results for Simple Insertion, Simple Deletion, and CAAP on ResNet-18, ResNet-50, DenseNet-161, and VGG-16. Deletion ($\downarrow$), Insertion ($\uparrow$), and Ins-Del ($\uparrow$) are reported. Best is bold, second best is underlined in each column.
@@ B.3 CNN architectures
Simple Insertion 16.05±0.67 38.75±1.20 22.70±1.13 15.65±0.64 49.11±1.28 33.46±1.25 20.83±0.78 52.87±1.34 32.04±1.26 14.60±0.62 41.01±1.22 26.41±1.14
Simple Deletion 12.54±0.57 54.44±1.19 41.90±1.09 14.74±0.78 62.19±1.18 47.45±1.22 20.35±0.95 65.98±1.16 45.64±1.17 11.66±0.60 57.54±1.17 45.88±1.09
@@
Table 15: Localization quantitative results for Simple Insertion, Simple Deletion, and CAAP on ResNet-18, ResNet-50, DenseNet-161, and VGG-16. AUPR1 ($\uparrow$), AUPR0 ($\uparrow$), and PG ($\uparrow$) are reported. Best is bold, second best is underlined in each column.
@@
Simple Insertion 36.91±2.23 65.50±2.37 47.66±4.41 42.00±2.21 69.03±2.28 56.25±4.38 39.70±2.24 67.83±2.19 49.22±4.42 35.12±2.15 64.95±2.31 42.19±4.37
Simple Deletion 50.49±2.20 72.09±2.31 60.94±4.31 49.40±2.08 69.96±2.35 64.06±4.24 49.54±2.14 71.87±2.25 62.50±4.28 51.61±2.23 73.62±2.32 64.84±4.22
@@
We further evaluate the intervention strategies on convolutional architectures, including ResNet-18, ResNet-50 [14], DenseNet-161 [15], and VGG-16 [32]. The results are reported in Tables 14 and 15. Compared with the transformer results reported earlier, the overall performance of intervention-based methods is lower across both faithfulness and localization metrics. In particular, the Insertion and Insertion-Deletion scores are substantially smaller than those observed for ViT models. One possible explanation is that transformer architectures are generally more robust to distribution shifts than convolutional networks. Since intervention-based attribution methods modify the input or intermediate representations, they introduce out-of-distribution (OOD) patterns. Transformers appear to handle these perturbations more reliably, which makes intervention-based attribution more suitable for transformer-based models.
@@
Within the CNN architectures, CAAP still achieves the best overall results in most cases. As shown in Table 14, CAAP produces the highest Insertion-Deletion scores across all four models and often improves the Insertion score, while maintaining competitive Deletion AUC. Localization metrics in Table 15 show the same trend, where CAAP consistently achieves the highest AUPR1 and AUPR0 and improves PG in most cases. Another difference from the transformer experiments is that the Simple Deletion baseline generally outperforms Simple Insertion on CNNs. This behavior suggests that removing pixels may better reflect the model’s reliance on local features in convolutional networks. In contrast, inserting isolated patches into a blank image provides limited evidence for the classifier, possibly due to the OOD nature of the input.
@@
Figure 11: Visualization of attribution maps produced by different methods across various models. The target object is an Granny smith, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 12: Visualization of attribution maps produced by different methods across various models. The target object is an Black swan, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 13: Visualization of attribution maps produced by different methods across various models. The target object is an Cheetah, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 14: Visualization of attribution maps produced by different methods across various models. The target object is a Sandbar, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 15: Visualization of attribution maps produced by different methods across various models. The target object is a Black grouse, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 16: Visualization of attribution maps produced by different methods across various models. The target object is a Goldfish, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 17: Visualization of attribution maps produced by different methods across various models. The target object is a Gyromitra, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 18: Visualization of attribution maps produced by different methods across various models. The target object is a Mousetrap, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 19: Visualization of attribution maps produced by different methods across various models. The target object is an Ear, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 20: Visualization of attribution maps produced by different methods across various models. The target object is a European fire salamander, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 21: Visualization of attribution maps produced by different methods across various models. The target object is a Tiger beetle, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 22: Visualization of attribution maps produced by different methods across various models. The target object is a Proboscis monkey, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 23: Visualization of attribution maps produced by different methods across various models. The target object is a White wolf, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 24: Visualization of attribution maps produced by different methods across various models. The target object is a Catamaran, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 25: Visualization of attribution maps produced by different methods across various models. The target object is a Lens cap, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 26: Visualization of attribution maps produced by different methods across various models. The target object is an Indigo bunting, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 27: Visualization of attribution maps produced by different methods across various models. The target object is a Cabbage butterfly, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 28: Visualization of attribution maps produced by different methods across various models. The target object is an Electric guitar, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 29: Visualization of attribution maps produced by different methods across various models. The target object is an Echidna, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 30: Visualization of attribution maps produced by different methods across various models. The target object is a Neck brace, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 31: Visualization of attribution maps produced by different methods across various models. The target object is an American alligator, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 32: Visualization of attribution maps produced by different methods across various models. The target object is an Eft, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 33: Visualization of attribution maps produced by different methods across various models. The target object is an American lobster, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 34: Visualization of attribution maps produced by different methods across various models. The target object is a Spotted salamander, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 35: Visualization of attribution maps produced by different methods across various models. The target object is a Cougar, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 36: Visualization of attribution maps produced by different methods across various models. The target object is a Lawn mower, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 37: Visualization of attribution maps produced by different methods across various models. The target object is an Albatross, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 38: Visualization of attribution maps produced by different methods across various models. The target object is a Trombone, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 39: Visualization of attribution maps produced by different methods across various models. The target object is a Junco, and our method more accurately delineates the object boundaries compared to the baselines.
@@
Figure 40: Visualization of attribution maps produced by different methods across various models. The target object is a Hermit crab, and our method more accurately delineates the object boundaries compared to the baselines.

### 17 May 2026 v2

@@ D.2 Input-Based Intervention Baselines
Input Insertion 55.63±1.24 69.51±1.19 10.63±1.01 43.86±1.10 75.33±0.87 14.93±0.83 49.79±1.20 79.41±1.11 19.59±0.79 33.33±0.92 64.50±0.97 22.16±0.75
Input Deletion 58.89±1.37 69.75±1.07 10.86±1.20 60.40±1.27 72.38±0.89 11.98±1.14 59.82±1.33 80.18±1.05 20.36±0.93 42.34±0.90 61.13±0.89 18.79±0.88
@@
Table 14: Localization quantitative results for Input Insertion, Input Deletion, and CAAP on ImageNet. AUPR1 ($\uparrow$), AUPR0 ($\uparrow$) and PG ($\uparrow$) are reported. Best is bold, second best is underlined in each column.
@@
Input Insertion 49.28±1.99 71.10±2.18 61.72±4.30 55.32±1.96 75.55±2.09 59.38±4.34 48.13±2.02 72.81±2.16 52.34±4.41 55.59±1.97 79.02±1.99 62.50±4.28
Input Deletion 40.82±2.07 67.53±2.23 46.09±4.41 40.44±2.10 66.01±2.18 58.59±4.35 39.05±2.09 67.45±2.27 45.31±4.40 47.51±2.50 80.74±1.69 57.81±4.37
@@
To better understand the role of representation-level interventions, we compare CAAP with two simple input-based alternatives: Input Insertion and Input Deletion. In Input Insertion, a selected source patch is inserted directly into a blank target image and the resulting class score is used as the attribution signal. In Input Deletion, the selected patch is removed from the input image while all other pixels remain unchanged. These procedures operate only at the input level and do not modify internal representations across layers. All methods use the default padding ($3\times 3$ patch region). Tables 13 and 14 report the results on ImageNet across four ViT backbones.
@@
D.3 CNN Architectures
Table 15: Faithfulness quantitative results for Input Insertion, Input Deletion, and CAAP on ResNet-18, ResNet-50, DenseNet-161, and VGG-16. Deletion ($\downarrow$), Insertion ($\uparrow$), and Ins-Del ($\uparrow$) are reported. Best is bold, second best is underlined in each column.
@@ D.3 CNN Architectures
Input Insertion 16.05±0.67 38.75±1.20 22.70±1.13 15.65±0.64 49.11±1.28 33.46±1.25 20.83±0.78 52.87±1.34 32.04±1.26 14.60±0.62 41.01±1.22 26.41±1.14
Input Deletion 12.54±0.57 54.44±1.19 41.90±1.09 14.74±0.78 62.19±1.18 47.45±1.22 20.35±0.95 65.98±1.16 45.64±1.17 11.66±0.60 57.54±1.17 45.88±1.09
@@
Table 16: Localization quantitative results for Input Insertion, Input Deletion, and CAAP on ResNet-18, ResNet-50, DenseNet-161, and VGG-16. AUPR1 ($\uparrow$), AUPR0 ($\uparrow$), and PG ($\uparrow$) are reported. Best is bold, second best is underlined in each column.
@@
Input Insertion 36.91±2.23 65.50±2.37 47.66±4.41 42.00±2.21 69.03±2.28 56.25±4.38 39.70±2.24 67.83±2.19 49.22±4.42 35.12±2.15 64.95±2.31 42.19±4.37
Input Deletion 50.49±2.20 72.09±2.31 60.94±4.31 49.40±2.08 69.96±2.35 64.06±4.24 49.54±2.14 71.87±2.25 62.50±4.28 51.61±2.23 73.62±2.32 64.84±4.22
@@
We further evaluate the intervention strategies on convolutional architectures, including ResNet-18, ResNet-50 [15], DenseNet-161 [16], and VGG-16 [33]. The results are reported in Tables 15 and 16. Compared with the transformer results reported earlier, the overall performance of intervention-based methods is lower across both faithfulness and localization metrics. In particular, the Insertion and Insertion-Deletion scores are substantially smaller than those observed for ViT models. One possible explanation is that transformer architectures are generally more robust to distribution shifts than convolutional networks. Since intervention-based attribution methods modify the input or intermediate representations, they introduce out-of-distribution (OOD) patterns. Transformers appear to handle these perturbations more reliably, which makes intervention-based attribution more suitable for transformer-based models.
@@
Within the CNN architectures, CAAP still achieves the best overall results in most cases. As shown in Table 15, CAAP produces the highest Insertion-Deletion scores across all four models and often improves the Insertion score, while maintaining competitive Deletion AUC. Localization metrics in Table 16 show the same trend, where CAAP consistently achieves the highest AUPR1 and AUPR0 and improves PG in most cases. Another difference from the transformer experiments is that the Input Deletion baseline generally outperforms Input Insertion on CNNs. This behavior suggests that removing pixels may better reflect the model’s reliance on local features in convolutional networks. In contrast, inserting isolated patches into a blank image provides limited evidence for the classifier, possibly due to the OOD nature of the input.
Appendix E Broader Impact
This work contributes to interpretability and attribution methods for Vision Transformers by proposing a causal intervention framework based on activation patching. Improved attribution quality may help researchers better understand model behavior, identify reliance on spurious correlations, and support debugging and auditing of vision systems in sensitive applications. At the same time, attribution methods may be over-interpreted as fully faithful explanations of model reasoning, despite inherent limitations and architecture-specific assumptions. Incorrect or overly confident interpretations could lead to misplaced trust in deployed systems. Our method is designed as a diagnostic and research tool rather than a guarantee of model transparency or safety.
All experiments use publicly available datasets and pretrained models, including ImageNet-1K, ImageNet-S, Oxford-IIIT-Pet, Food-101, CLIP, DINOv2, DeiT3, and ViT models, under their respective research licenses and terms of use. The paper introduces a new attribution method but does not release new datasets or generative models.
Appendix F Experimental Sampling and Resources
For each dataset, we evaluate all methods on a fixed random subset of 500 samples. The same samples are used across baselines and ablations to ensure fair comparison. When segmentation annotations are required, samples are drawn only from images with valid object-level annotations. All experiments were run on four NVIDIA RTX 4090 GPUs, with CPU resources used for preprocessing, metric computation, and result aggregation.
@@ Appendix F Experimental Sampling and Resources
Figure 13: Visualization of single-object attribution maps produced by different methods across various models. The target object is an Granny smith.
@@
Figure 14: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Black swan.
@@
Figure 15: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Cheetah.
@@
Figure 16: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Sandbar.
@@
Figure 17: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Black grouse.
@@
Figure 18: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Goldfish.
@@
Figure 19: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Gyromitra.
@@
Figure 20: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Mousetrap.
@@
Figure 21: Visualization of single-object attribution maps produced by different methods across various models. The target object is an Ear.
@@
Figure 22: Visualization of single-object attribution maps produced by different methods across various models. The target object is a European fire salamander.
@@
Figure 23: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Tiger beetle.
@@
Figure 24: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Proboscis monkey.
@@
Figure 25: Visualization of single-object attribution maps produced by different methods across various models. The target object is a White wolf.
@@
Figure 26: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Catamaran.
@@
Figure 27: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Lens cap.
@@
Figure 28: Visualization of single-object attribution maps produced by different methods across various models. The target object is an Indigo bunting.
@@
Figure 29: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Cabbage butterfly.
@@
Figure 30: Visualization of single-object attribution maps produced by different methods across various models. The target object is an Electric guitar.
@@
Figure 31: Visualization of single-object attribution maps produced by different methods across various models. The target object is an Echidna.
@@
Figure 32: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Neck brace.
@@
Figure 33: Visualization of single-object attribution maps produced by different methods across various models. The target object is an American alligator.
@@
Figure 34: Visualization of single-object attribution maps produced by different methods across various models. The target object is an Eft.
@@
Figure 35: Visualization of single-object attribution maps produced by different methods across various models. The target object is an American lobster.
@@
Figure 36: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Spotted salamander.
@@
Figure 37: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Cougar.
@@
Figure 38: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Lawn mower.
@@
Figure 39: Visualization of single-object attribution maps produced by different methods across various models. The target object is an Albatross.
@@
Figure 40: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Trombone.
@@
Figure 41: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Junco.
@@
Figure 42: Visualization of single-object attribution maps produced by different methods across various models. The target object is a Hermit crab.
Figure 43: Visualization of multi-object attribution maps produced by a selected set of strong methods. The target objects are a dolphin and a shark.
Figure 44: Visualization of multi-object attribution maps produced by a selected set of strong methods. The target objects are a pink flower and a blue butterfly.
Figure 45: Visualization of multi-object attribution maps produced by a selected set of strong methods. The target objects are a globe, a bed, and books.
Figure 46: Visualization of multi-object attribution maps produced by a selected set of strong methods. The target objects are a gnome, a unicycle, a clock, and a duck crossing sign.
Figure 47: Visualization of multi-object attribution maps produced by a selected set of strong methods. The target objects are a dog, shoes, and a camera.
Figure 48: Visualization of multi-object attribution maps produced by a selected set of strong methods. The target objects are a playground slide, a ponytail holder, and a red bag.
Figure 49: Visualization of multi-object attribution maps produced by a selected set of strong methods. The target objects are a panda, a clownfish, and hearts.
Figure 50: Visualization of multi-object attribution maps produced by a selected set of strong methods. The target objects are shoes and purple socks.
Figure 51: Visualization of multi-object attribution maps produced by a selected set of strong methods. The target objects are a laptop, a tray of food, and pillows.
Figure 52: Visualization of multi-object attribution maps produced by a selected set of strong methods. The target objects are a cat, an umbrella, a girl, and eyes.